OpenStudy (anonymous):

multiply and simplify the expression (5vv-3)(6v-5)=

OpenStudy (anonymous):

95V-3)(6V-5)=

OpenStudy (anonymous):

(5v-3)(6v-5)=

OpenStudy (anonymous):

Are you trying to find v? Or, just the expanded version?

OpenStudy (anonymous):

i'm trying to simplify the expression

OpenStudy (anonymous):

(5v-3)(6v-5) Multiply each term in the first group by each term in the second group using the FOIL method. FOIL stands for First Outer Inner Last, and is a method of multiplying two binomials. First, multiply the first two terms in each binomial group. Next, multiply the outer terms in each group, followed by the inner terms. Finally, multiply the last two terms in each group. (5v*6v+5v*-5-3*6v-3*-5) Multiply 5v by 6v to get 30v^(2). (30v^(2)+5v*-5-3*6v-3*-5) Multiply 5v by -5 to get -25v. (30v^(2)-25v-3*6v-3*-5) Multiply -3 by 6v to get -18v. (30v^(2)-25v-18v-3*-5) Multiply -3 by -5 to get 15. (30v^(2)-25v-18v+15) According to the distributive property, for any numbers a, b, and c, a(b+c)=ab+ac and (b+c)a=ba+ca. Here, v is a factor of both -25v and -18v. (30v^(2)+(-25-18)v+15) Subtract 18 from -25 to get -43. (30v^(2)+(-43)v+15) Remove the parentheses. (30v^(2)-43v+15) Remove the parentheses around the expression and you get: 30v^(2)-43v+15